OpenStudy (aaronandyson):

10g of ice at 0 degrees absorbs 5460 J of heat energy to melt and change to water at 50 degrees C. Calculate the specific latent heat of fusion of ice.Specific heat capacity of water is 4200 J kg^-1 K^-1

OpenStudy (aaronandyson):

@Michele_Laino

OpenStudy (michele_laino):

the requested amount of heat \(Q\) to go from ice at \(0\) degrees to water at \(50\) degrees, is: \[\huge Q = Lm + cm\Delta \theta \] where \(m=0.01\)Kg, \(L\) is the latent heat, \(\Delta \theta=50\), \(c=4200\)

OpenStudy (michele_laino):

of course \(Q=5460\)

OpenStudy (michele_laino):

remark: (\( \Delta \theta= 50 °C= 50 °K\))

OpenStudy (aaronandyson):

5460 = 0.04 times L times 2100??

OpenStudy (aaronandyson):

5460 - 0.04 times L plus 2100*

OpenStudy (michele_laino):

no, please we have: \(5460= L m+2100\)

OpenStudy (michele_laino):

so: \[\huge L = \frac{{5460 - 2100}}{{0.010}} = ...?\]

OpenStudy (aaronandyson):

336000

OpenStudy (michele_laino):

correct!

OpenStudy (aaronandyson):

whats the unit..??

OpenStudy (michele_laino):

please don't forget the right unit of measure It is \(Joules/Kg\)